OpenStudy (brettross25):

If f(x) = 2x + 3 and g(x) = x^2 − 1, find the values of combining these functions. Match each combined function to its corresponding value. (f + g)(2) (f − g)(4) (f ÷ g)(2) (f × g)(1) .............................. -4 10 0 7/3

OpenStudy (zzr0ck3r):

\((f+g)(2)=f(2)+g(2)=(2*2+3)+(2^2-1)\)

OpenStudy (zzr0ck3r):

Do something similar for the other ones.

OpenStudy (zzr0ck3r):

Does this makes sense @BrettRoss25 ?

OpenStudy (brettross25):

idk

OpenStudy (brettross25):

i dont get it

OpenStudy (prathamesh_m):

(f+g)(2) = f(2) + g(2) Which means you simply substitute 2 in f(x) and g(x) to find f(2) and g(2). Then you add them to get (f+g)(2)

OpenStudy (prathamesh_m):

Similarly for (f x g)(2), you multiply instead of add and so on.

OpenStudy (skullpatrol):

The value of a function f at 2 is denoted by f(2). Any value of a function can be found by replacing the variable in the defining equation of the function by the given number. So replace the variable in your functions by 2 and then perform the operations asked for in the question.